SOHO DAILY REPORT Operations Day - 169/2003 (18 June 03) Mission Day 2756 SOHO FLIGHT OPERATIONS TEAM, NASA/GSFC Ref. #: SH-FOT-DR-030618 ==================================================================== PASS #1 Madrid DSS66 0413-1336 PASS #2 Goldstone DSS16 1316-2117 PASS #3 Canberra DSS46 2104-170/0355 ********************************************************************* PASS #1 Madrid DSS66 0413-1336 Anomalies: G03-0438 0414 POCC timing problem caused by intermittent telemetry while executing the pre-pass limit set procedure. CDS limits were inhibited from 0414 until 1731; no other impacts. NRT: 1156-1244 Activities Summary: - SSR Dump - Collected MDI-M (0637-0653) - Collected MDI-H (0653-1306, 1314-1336) - OBT Distribution to SWAN, CELIAS - Collected Ranging Data - SSR SEF Reset (x2) ------------------ PASS #2 Goldstone DSS16 1316-2117 Anomalies: G03-0436 1413 Lost 7 minutes of unrecoverable TLM from 1413 to 1420, 6 minutes of recoverable TLM from 1613 to 1619, and 45 seconds of unrecoverable TLM at 1813 all due to a MPA problem with the predicts at D16. DR #G102917 G03-0437 1513 POCC timing problem while executing procedure to set ESR warning flag receivers on board the spacecraft. NRT: 1408-1413 1423-1449 Activities Summary: - SSR Dump - Collected MDI-M (1350-1502, 1745-1917) - Collected MDI-H (1316-1350) - Uplinked delayed load SWA030617131 - Reaction Wheel Speed Monitoring Daily Adjustment - CRS Drift Estimation Update - SSR SEF Reset (x2) - SSR Record (x2) - Executed CS-14: Loss of TM in High Rate - Executed Nominal Ops Script: Profile in NM/RMW - Executed OCD #1787: EOF Inputs for Off-point Activities - Executed OCD #1788: HGA Test Maneuver - Executed OCD #1789: Switch to APME-B and Repoint HGA ------------------ PASS #3 Canberra DSS46 2104-170/0355 Anomalies: None NRT: 2319-2326 0008-0133 Activities Summary: - SSR Dump - Collected MDI-M (0007-0335) - OBT Distribution to SWAN, CELIAS - Collected Ranging Data - SSR SEF Reset (x2) - SSR Record - Executed OCD #1789: Switch to APME-B and Repoint HGA - Executed OCD #1790: Raise HGA Z-Motor Heater Temperature ------------------ -- End of Report --